OpenStudy (anonymous):

If (x^2/16) + (y^2/49) = 1 and y(2)=(7/4)sqrt12 Find y'(2) by implicit differentiation

OpenStudy (anonymous):

Please help I am very confused and lost as to what to do with this problem.

OpenStudy (callisto):

Do you know how to find y' ?

OpenStudy (anonymous):

I don't know I don't know if this is correct but this is what I started doing I tried to find the derivative of the one with the x and y's and got (32x/16^2)+ (98y/49^2)(dy/dx)

OpenStudy (callisto):

Hmm.. \[\frac{x^2}{16} + \frac{y^2}{49} = 1\]Diff. both sides w.r.t. x. How do you differentiate the first term on the left?

OpenStudy (anonymous):

I tried to do the bottom multiplied by the derivative of the top - top multiplied by the derivative of the bottom over the bottom squared for the left and then the derivative of 1 just goes to 0

OpenStudy (callisto):

Hmm.. So, we agree that derivative of one is zero. No problem. The problem is finding the derivative of the term on the left. \[\frac{d}{dx}(\frac{x^2}{16} + \frac{y^2}{49})=\frac{d}{dx}(\frac{x^2}{16}) + \frac{d}{dx}(\frac{y^2}{49})\]So far so good?

OpenStudy (anonymous):

Yes

OpenStudy (callisto):

Nice~ Consider the first one \[\frac{d}{dx}(\frac{x^2}{16}) = \frac{1}{16}(\frac{d}{dx}x^2)\]Happy with this?

OpenStudy (anonymous):

Okay so then y would be 1/49((d/dx)y^2) correct?

OpenStudy (callisto):

Oh yes! Can you find the derivatives again?

OpenStudy (anonymous):

So do you mean by the inside which would be d/dx 2x and d/dx 2y?

OpenStudy (callisto):

Hmm, not really correct. What is the derivative of x^2 with respect to x?

OpenStudy (anonymous):

2x

OpenStudy (callisto):

Yes. So, \(\frac{1}{16}(\frac{d}{dx}x^2) = \frac{1}{16}(2x)\), right?

OpenStudy (anonymous):

yes so then 1/49(2y) is the other part correct?

OpenStudy (callisto):

1/49(2y) => not correct.

OpenStudy (callisto):

Imagine this.. \[\frac{d}{dx}x^2 = 2x\frac{dx}{dx} = 2x\]\[\frac{d}{dx}y^2 = ...?\]

OpenStudy (anonymous):

2y then would follow that logic correct?

OpenStudy (callisto):

Nope, not really, instead, it should be (2y)y'